# Radar Perception in Autonomous Driving: Exploring Different Data Representations

### v1

@@ IV-A1 Consideration of Application Scenarios
For applications requiring accurate object detection and tracking, representations like radar tensors or point clouds prove valuable. Radar tensors, such as RAD tensors, offer visualizations of range, Doppler velocity, and azimuth, which can be used to extract features using image-based methods [47, 97]. Point cloud representations deliver 3D spatial information, aligned with real-world object locations, enabling utilization with point-based network architectures [154, 155]. Additionally, point clouds are the output format of conventional radar, benefiting from well-developed product suites and research materials [156].
@@ IV-A2 Consideration of Computational Efficiency
IV-B How to Improve the Performance of Radar Perception?
The constraints inherent in radar sensors and the limitations associated with radar data processing algorithms present challenges to the performance of radar perception. Improving radar perception performance is crucial for autonomous driving, enabling more accurate decisions and predictions. In the following, we explore four areas: creating high-quality datasets, dealing with inherent limitations, improving data processing algorithms, and employing 4D radar sensors.
@@ IV-B How to Improve the Performance of Radar Perception?
Creating High-Quality Datasets
Calibration Annotation Develop adaptive and self-calibration algorithms Design real-time and dynamic calibration techniques Explore auto-labeling techniques
Dealing with Inherent Limitations
Sparsity and limited resolution Noise and clutter Explore the distribution of radar output Construct features based on physical properties Model inherent uncertainty
Improving Data Processing Algorithms
Parameter estimation Feature extraction Use neural networks instead of traditional FFTs Apply graph neural network to extract features Utilize occupancy prediction for scene understanding
Employing 4D Radar Sensors
Large size Perception algorithms Extract core values from tensor dimensions Develop efficient 3D object perception models Design specific architectures on dense information
@@ IV-B4 Employing 4D Radar Sensors
Radar sensors have made significant progress, transitioning from 3D to 4D capabilities, including enhanced resolution and elevation measurement capabilities. As a result, more and more research is focusing on the 4D radar, with growing numbers of radar datasets and algorithms proposed recently. Given these advancements, employing 4D radar sensors is a promising and potential research direction for radar perception in the broader domain of autonomous driving. However, 4D radar is still challenging in terms of large data size as well as perception algorithms.
@@
IV-C What about Data Fusion for Radar Representation?
While radar sensors are good at detecting and ranging objects, they provide limited information about object characteristics like color, texture, and fine-grained visual details. Integrating radars with complementary sensors, such as cameras or LiDARs, overcomes this limitation and provides comprehensive perception solutions [1, 208]. However, radar data fusion also faces some challenges, particularly in heterogeneous data association and in challenging condition processing.
@@ IV-C1 Associating Heterogeneous Data
A significant challenge is the ambiguity in associating radar data with other modalities, as they are heterogeneous. For radar-camera fusion, a conventional approach involves projecting radar data onto the image plane and subsequently establishing their correspondence via a calibration matrix [209, 101, 168, 210]. For radar-LiDAR fusion, one common approach exploits geometric information to associate radar and LiDAR point clouds, which compares the position, velocity, and shape of the detected objects to determine if they correspond to the same physical object [211]. However, projection-based or geometric-based association results in poor alignment with object centers. As previously stated, radar data exhibits sparsity, inaccuracy, and noise, thereby causing inadequate associations in both object-level and data-level fusion scenarios.
Therefore, associating radar data with other modalities is a critical but challenging question. In our opinion, the incorporation of BEV features, transformer architectures, and attention mechanisms are valuable research directions that can significantly improve fusion performance. BEV provides a top-down perspective view of the vehicle’s surroundings, thereby enhancing spatial understanding. Using BEV transformations and transformer architectures, radar-camera or radar-LiDAR fusion methods (e.g., RCBEV [212], CRAFT [213], CRN [214], RCFusion [215], LXL [123], Bi-LRFusion [216], BEVGuide [217]) have proven to deliver impressive performance.
Moreover, attention mechanisms play a crucial role in determining the relevance of different sensor modalities at different spatial or temporal locations. In the case of data association, the attention mechanism filters occluded radar returns and mitigates background clutter, which facilitates the alignment and matching of objects or features across the different modalities [170, 213]. Thus, attention-based association with adaptive thresholds emerges as a prospective approach for establishing a connection between radar and other modalities. Furthermore, the Joint Probabilistic Data Association (JPDA) method is a potential research direction that leverages Bayesian filtering techniques to estimate the likelihood of associations based on the joint probability distribution of radar and other modalities [218]. By considering uncertainties associated with the measurements and estimates, JDPA enables robust data association.
@@ IV-C2 Handling Challenging Conditions
Fusion in challenging conditions refers to the ability of a sensor fusion system to maintain reliable and accurate perception even in adverse environments, where individual sensors may face difficulties. In such conditions, sensor measurements are noisy, degraded, or influenced by poor visibility, adverse weather and object occlusions [219, 220, 221]. Therefore, sensor fusion aims to leverage the complementary strengths of different sensors to compensate for the limitations of one sensor with the capabilities of another. Developing robust fusion algorithms that handle varying environmental conditions and adaptively combine sensor data is critical for reliable perception.
Fusion in challenging conditions may require adaptive fusion strategies that dynamically adjust fusion parameters based on environmental conditions. Developing data weighting techniques that assign appropriate weights to sensor measurements based on reliability can enhance fusion performance under challenging conditions [222, 223]. For example, under adverse weather conditions, radar sensors may be more reliable than cameras or LiDAR, and therefore the fusion algorithm should adaptively emphasize radar data and de-emphasize other sensors.
Effective fusion in challenging conditions also requires uncertainty estimation associated with sensor measurements. Probabilistic fusion techniques should be incorporated to model and propagate uncertainties properly. Bayesian neural networks are notable techniques employed for uncertainty estimation, using the prior distribution of network weights to infer the posterior distribution. This enables the computation of probabilities for specific predictions, thereby enhancing the reliability of the fusion network [208, 224]. Such techniques should enable fusion algorithms to make informed decisions considering the uncertainty levels in fused data.
@@ V Conclusion
Through analyzing various radar data representations, we thoroughly discuss the key challenges and propose potential research directions for radar perception. In general, radar perception is progressing towards data representations containing rich information. On the one hand, representations like ADC signals and radar tensors offer increased potential information, thereby holding significant value for radar perception. On the other hand, new 4D radar sensors bring forth denser point clouds and higher resolutions, representing a noteworthy trend in autonomous driving. Moreover, emerging radar-based sensor fusion works are also hot topics in autonomous driving and are expected in future works. Above all, we hope our review serves as a valuable reference for both researchers and practitioners in developing robust radar perception, making our vehicles and transportation systems safer and more efficient.
@@ Acknowledgment
This research was funded by Suzhou Municipal Key Laboratory for Intelligent Virtual Engineering (SZS2022004), Research Development Fund of XJTLU (RDF-19-02-23), XJTLU AI University Research Centre, Jiangsu Province Engineering Research Centre of Data Science and Cognitive Computation at XJTLU, SIP AI innovation platform (YZCXPT2022103), Suzhou Science and Technology Project (SYG202122) and Key Programme Special Fund of Xi’an Jiaotong-Liverpool University (KSF-A-19). This work received financial support from Jiangsu Industrial Technology Research Institute (JITRI) and Wuxi National Hi-Tech District (WND).

### v2

@@ IV-A1 Consideration of Application Scenarios
For applications requiring accurate object detection and tracking, representations like radar tensors or point clouds prove valuable. Radar tensors, such as RAD tensors, offer visualizations of range, Doppler velocity, and azimuth, which can be used to extract features using image-based methods [47, 103]. Point cloud representations deliver 3D spatial information aligned with real-world object locations, enabling utilization with point-based network architectures [167, 168]. Additionally, point clouds are the output format of conventional radar, benefiting from well-developed product suites and research materials [169].
@@ IV-A2 Consideration of Computational Efficiency
IV-B How to Improve Radar Perception Performance?
The constraints inherent in radar sensors and the limitations associated with radar data processing algorithms present challenges to the performance of radar perception. Improving radar perception performance is crucial for autonomous driving, enabling more accurate decisions and predictions. In the following, we explore four areas: creating high-quality datasets, dealing with inherent limitations, improving data processing algorithms, and employing 4D radar sensors. An overview of challenges and research directions in these four areas is described in Table IV.
@@ IV-B How to Improve Radar Perception Performance?
Creating High-Quality Datasets Calibration Annotation Develop adaptive and self-calibration algorithms Design real-time and dynamic calibration techniques Explore auto-labeling techniques
Dealing with Inherent Limitations Sparsity and limited resolution Noise and clutter Explore the distribution of radar output Construct features based on physical properties Model inherent uncertainty
Improving Data Processing Algorithms Parameter estimation Feature extraction Use neural networks instead of traditional FFTs Apply graph neural network to extract features Utilize occupancy prediction for scene understanding
Employing 4D Radar Sensors Large size Perception algorithms Extract core values from tensor dimensions Develop efficient 3D object perception models Design specific architectures on dense information
@@ IV-B4 Employing 4D Radar Sensors
Radar sensors have made significant progress, transitioning from 3D to 4D capabilities, including enhanced resolution and elevation measurement capabilities. As a result, more and more research is focused on the 4D radar, with growing numbers of radar datasets and algorithms proposed recently. Given these advancements, employing 4D radar sensors is a promising and potential research direction for radar perception in the broader domain of autonomous driving. However, 4D radar is still challenging in terms of large data size as well as perception algorithms.
@@
IV-C What about Data Fusion for Radar Representations?
While radar sensors are good at detecting and ranging objects, they provide limited information about object characteristics like color, texture, and fine-grained visual details. Integrating radars with complementary sensors, such as cameras or LiDARs, overcomes this limitation and provides comprehensive perception solutions [1, 221]. However, radar data fusion also faces some challenges, particularly in heterogeneous data association and in challenging condition processing. An overview of these challenges and research directions is summarized in Table V.
TABLE V: Overview of challenges and research directions on radar data fusion.
Topic Challenges Research Directions
Associating Heterogeneous Data Projection-based or geometric-based associations result in ineffective alignment Sensitive to occluded radar returns and background clutter Utilize BEV features to enhance spatial understanding Apply attention mechanisms to mitigate background clutter Employ joint probabilistic data association methods to estimate the likelihood of associations
Handling Challenging Conditions Sensitive to adverse weather and object occlusions Uncertainty estimation of different sensor measurements Develop adaptive fusion strategies to adjust fusion parameters based on environmental conditions Employ probabilistic fusion techniques to model and propagate uncertainties
@@ IV-C1 Associating Heterogeneous Data
A significant challenge is the ambiguity in associating radar data with other modalities, as they are heterogeneous. For radar-camera fusion, existing approaches involve projecting radar data onto the image plane and subsequently establishing their correspondence via a calibration matrix [222, 108, 181, 223]. For radar-LiDAR fusion, existing methods exploit geometric information to associate radar and LiDAR point clouds by comparing the position, velocity, and shape of the detected objects to determine if they correspond to the same physical object [224]. However, projection-based or geometric-based associations result in ineffective alignment with object centers. As previously stated, radar data exhibits sparsity, inaccuracy, and noise, thereby causing inadequate associations in both object-level and data-level fusion scenarios.
Therefore, associating radar data with other modalities is a critical but challenging question. In our opinion, the incorporation of BEV features, transformer architectures, and attention mechanisms are valuable research directions that can significantly improve fusion performance. BEV provides a top-down perspective view of the vehicle’s surroundings, thereby enhancing spatial understanding. Using BEV transformations and transformer architectures, radar-camera fusion methods (e.g., RCBEV [225], CRAFT [226], CRN [227], RCFusion [228], LXL [135]) or radar-LiDAR fusion methods (e.g., Bi-LRFusion [229], BEVGuide [230], ACF-Net [231], ST-MVDNet++ [232]) have proven to deliver impressive performance.
Moreover, attention mechanisms have been successfully utilized to deal with complex contextual relationships, which makes them promising for data fusion tasks [1]. Attention mechanisms play a crucial role in determining the relevance of different sensor modalities at different spatial or temporal locations. In the case of data association, the attention mechanism filters occluded radar returns and mitigates background clutter, which facilitates the alignment and matching of objects or features across the different modalities [183, 226]. Thus, attention-based association with adaptive thresholds emerges as a prospective approach for establishing a connection between radar and other modalities. Furthermore, the Joint Probabilistic Data Association (JPDA) method is a potential research direction that leverages Bayesian filtering techniques to estimate the likelihood of associations based on the joint probability distribution of radar and other modalities [233]. By considering uncertainties associated with the measurements and estimates, JDPA enables robust data association.
@@ IV-C2 Handling Challenging Conditions
Fusion in challenging conditions refers to the ability of a sensor fusion system to maintain reliable and accurate perception even in adverse environments, where individual sensors may face difficulties. In such conditions, sensor measurements are noisy, degraded, or influenced by poor visibility, adverse weather and object occlusions [234, 235, 236, 237]. Therefore, sensor fusion aims to leverage the complementary strengths of different sensors to compensate for the limitations of one sensor with the capabilities of another. Developing robust fusion algorithms that handle varying environmental conditions and adaptively combine sensor data is critical for reliable perception.
Fusion in challenging conditions may require adaptive fusion strategies that dynamically adjust fusion parameters based on environmental conditions. Developing data weighting techniques that assign appropriate weights to sensor measurements based on reliability can enhance fusion performance under challenging conditions [238, 239]. For example, radar sensors may be more reliable than cameras and LiDARs under adverse weather conditions. Therefore, the fusion algorithm should adaptively emphasize radar data and de-emphasize other sensors.
Effective fusion in challenging conditions also requires uncertainty estimation associated with sensor measurements. It refers to the lack of complete confidence or knowledge about the observations obtained from different sensors in the fusion process. Probabilistic fusion techniques should be incorporated to model and propagate uncertainties properly. Bayesian neural networks are notable techniques employed for uncertainty estimation, using the prior distribution of network weights to infer the posterior distribution. This enables the computation of probabilities for specific predictions, thereby enhancing the reliability of the fusion network [221, 240]. Such techniques should enable fusion algorithms to make informed decisions considering the uncertainty levels in fused data.
@@ V Conclusion
Through analyzing various radar data representations, we thoroughly discuss the key challenges and propose potential research directions for radar perception. In general, radar perception is progressing towards data representations containing rich information. On the one hand, representations like ADC signals and radar tensors offer increased potential information, thereby holding significant value for radar perception. On the other hand, new 4D radar sensors bring forth denser point clouds and higher resolutions, representing a noteworthy trend in autonomous driving. In terms of radar perception networks, the incorporation of transformer architectures, attention mechanisms, and BEV features provides valuable research directions to improve the perception performance. Moreover, emerging radar-based sensor fusion works are also hot topics in autonomous driving and are expected in future works. Above all, we hope our review serves as a valuable reference for both researchers and practitioners in developing robust radar perception, making our vehicles and transportation systems safer and more efficient.
@@ Acknowledgment
This research was funded by Suzhou Municipal Key Laboratory for Intelligent Virtual Engineering (SZS2022004), Research Development Fund of XJTLU (RDF-19-02-23), XJTLU AI University Research Centre, Jiangsu Province Engineering Research Centre of Data Science and Cognitive Computation at XJTLU, SIP AI innovation platform (YZCXPT2022103) and Suzhou Science and Technology Project (SYG202122). This work received financial support from Jiangsu Industrial Technology Research Institute (JITRI) and Wuxi National Hi-Tech District (WND).